OpenStudy (anonymous):

Let f(x) = 24/3x-2 Find f(-2). a. -6 b. -3 c. 3 d. 6

OpenStudy (unklerhaukus):

sub in x=-2

OpenStudy (chris911):

lol again which one dou think it is ??

OpenStudy (anonymous):

c

OpenStudy (chris911):

u r correct

OpenStudy (unklerhaukus):

f(x) = 24/(3x-2) f(-2) = 24/(3(-2)-2)

OpenStudy (unklerhaukus):

[it's not c]

OpenStudy (chris911):

how so this pic will show how I got it

OpenStudy (chris911):

it wont upload the pic but someone got the same answer as me http://openstudy.com/study#/updates/5214b63ae4b00ebc7dc021ce